OpenStudy (cutiekawdfi):

Find the focus and the directrix of the graph of the equation 1. y=5x^2

OpenStudy (agl202):

y = 5x^2, V(0,0) 5 = 1/4p, p = 1/20. y = -1/20 The directrix (Opening Up)

OpenStudy (cutiekawdfi):

is that it

OpenStudy (anonymous):

The graph will look like this: http://www.algebra.com/services/rendering/simplifier.mpl

OpenStudy (anonymous):

@cutiekawdfi